OpenStudy (anonymous):

The formula for electrical power is p=iV Which is the equivalent equation solved for i?

OpenStudy (anonymous):

I = P/V

OpenStudy (anonymous):

They have one like this p/v = I

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

The formula for the area of a triangle is a= 1/2 bh Which is the equivalent equation solved for h?

OpenStudy (anonymous):

h = 2a/b

OpenStudy (anonymous):

is it this one 2Ab=h

OpenStudy (anonymous):

\[\frac{ 2a }{ b }\]

OpenStudy (anonymous):

Which represents the equation 6x+3y=12 when solved for y?

OpenStudy (anonymous):

y = -2x + 4

OpenStudy (anonymous):

Leah is writing an equivalent equation which solves for y. What should be her next step? 3x+4y=8 3x-3x+4y=8-3x 4y=8-3x

OpenStudy (anonymous):

That's my last one I promise! Thanks so much for all the help! :)

OpenStudy (anonymous):

Are those the three options?

OpenStudy (anonymous):

I guess that was part of the question those aren't the awnsers tho

OpenStudy (anonymous):

I understand what they are asking now, the next step would be to divide both sides by 4

OpenStudy (anonymous):

Thanks! I made a 90!

OpenStudy (anonymous):

Congrats!